A triangle has an area of 77 square in. Find the length of the base if the base is 3 inches more then the height.
Found 2 solutions by tvandenberg, solver91311:
Answer by tvandenberg(45)

A = 1/2*B*H = 77 ---> B*H = 77*2 = 154
B = H + 3
(H+3)*H = 154
H^2 + 3H = 154
H^2 + 3H - 154 = 0
....quadratic eqn....
H = 11, -14
Can't be negative, H = 11
B = 11 + 3 = 14

Answer by solver91311(24713)


The area of a triangle is given by:



You are given that:



Substituting:







Solve the quadratic by factoring and exclude the extraneous negative root. The positive root will be the measure of the base of the triangle.

John